# Investigating and Mitigating Barren Plateaus in Variational Quantum Circuits: A Survey

### v1

@@ Abstract
In recent years, variational quantum circuits (VQCs) have been widely explored to advance quantum circuits against classic models on various domains, such as quantum chemistry and quantum machine learning. Similar to classic machine-learning models, VQCs can be optimized through gradient-based approaches. However, the gradient variance of VQCs may dramatically vanish as the number of qubits or layers increases. This issue, a.k.a. Barren Plateaus (BPs), seriously hinders the scaling of VQCs on large datasets. To mitigate the exponential gradient vanishing, extensive efforts have been devoted to tackling this issue through diverse strategies. In this survey, we conduct a systematic literature review of recent works from both investigation and mitigation perspectives. Besides, we propose a new taxonomy to categorize most existing mitigation strategies. At last, we provide insightful discussion for future directions of BPs.
@@
Variational Quantum Circuits, Barren Plateau, Survey
@@ I Introduction
In the era of noisy intermediate-scale quantum (NISQ), quantum computing has achieved significant advancement in various domains [1], such as quantum chemistry [2], quantum machine learning [3, 4, 5], and quantum architecture [6, 7]. Despite these advancements, the superiority of quantum computing over classical computing in the foreseeable future remains uncertain. One uncertainty arises during the optimization of Variational Quantum Circuits (VQCs), which have emerged as pivotal models for advancing quantum computing. Similar to classical models, VQCs can be optimized by gradient-based approaches [8, 9]. During optimization, however, the training of VQCs may be initially trapped in a flat landscape as the model size increases. McClean et al. [10] first categorize this issue as barren plateaus (BPs) in VQCs and ascertain that the variance of the gradient will exponentially decrease as the model size increases when the VQCs match the 2-design Haar distribution. Under this occurrence, most gradient-based techniques would fail. Thus, it is essential to mitigate the barren plateau issues in VQCs.
@@
  • We conduct comprehensive literature reviews on recent works about barren plateaus (BPs).
@@ II Preliminary of Barren Plateaus
$U(\theta)=U(\theta_{1},...,\theta_{L})=\prod_{l=1}^{L}U_{l}(\theta_{l})W_{l},$ (1)
where $U_{l}(\theta_{l})=e^{-i\theta_{l}V_{l}}$, $V_{l}$ is a Hermitian operator, and $W_{l}$ is unitary operator that doesn’t depend on $\theta_{l}$. $L$ denotes the number of layers in VQCs.
Similar to classical models, VQCs can be optimized by gradient-based methods. The cost function for optimization can be denoted as $C(\theta)$. Thus, $C(\theta)$ can be formulated as the expectation over Hermitian operator $H$:
@@
Based on the above loss function $C(\theta)$, we can further denote its gradient as $\frac{\partial C(\theta)}{\partial\theta_{l}}$, a.k.a. $\partial C$.
In 2018, McClean et al. [10] first investigated barren plateaus and revealed that under the assumption of the two-design Haar distribution, the variance of the gradient $Var[\partial C]$ in VQCs would exponentially decrease to zero during optimization as the model size, such as the number of qubits or layers, increases. To generally define the BPs, Qi et al. [11] provide a simple but generic definition as follows:
@@
where $F(N)\in o\left(\frac{1}{b^{N}}\right)$, for some $b>1$ and $N$ is the number of qubits in VQCs.
According to Equation 3, $Var[\partial C]$ will exponentially decrease to zero when the number of qubits $N$ increases. In this case, most gradient-based approaches will fail to train VQCs. Therefore, it is critical to circumvent the barren plateau issues for robustly training VQCs.
@@ III-A Understanding Barren Plateau Phenomena
To dive deeper into BPs, Wang et al. [12] further prove that the gradient will vanish exponentially under the consideration of local Pauli noise, which is quite different from the noise-free setting in [10]. Similarly, Singkanipa et al. [13] explore the effects of noise on VQCs, focusing on how non-unital noise contributes to the formation of barren plateaus and fixed points in the optimization landscape. Marrero et al. [14] show that excessive entanglement between the visible and hidden units in VQCs can hinder learning capacity. Holmes et al. [15] first investigate the scrambling processes and show that showing that any variational ansatz is highly probable to have a barren plateau landscape. Later, Holmes et al. [16] study the expressibility of VQCs and reveal that highly expressive VQCs may exhibit flatter optimization landscapes. Similarly, Ragone et al. [17] provide a unified framework to study the expressibility of VQCs. Liu et al. [18] theoretically provide an understanding of the distinction between BPs and “laziness”. In this study, laziness refers to the exponential suppression of variational angle updates during gradient descent. By studying laziness, researchers have a deeper understanding of the ansatz structure as a trade-off between flat landscapes and trainability, i.e., trainability in the presence of BPs via laziness). Friedrich et al. [19] examine how the dimensionality of qudits influences the occurrence of BPs in VQCs. Fontana et al. [20] introduce the concept of adjoint differentiation to analyze BPs and provide theoretical insights for designing quantum ansätze to avoid barren plateaus. Comparably, Diaz et al. [21] extend the theoretical understanding of BPs beyond the conventional dynamical Lie algebra (DLA) framework and propose new criteria to identify barren plateaus. Last, Anschuetz et al. [22] investigate other untrainable scenarios beyond barren plateaus.
@@ III-B Investigations from Model Perspectives
A portion of the research works aims to investigate BPs from the perspective of existing quantum circuits. Wiersema et al. [23] first examine a family of VQCs, named the Hamiltonian Variational Ansatz (HVA), and reveal that HVA is absent from BPs due to its favorable structural properties. Mao et al. [24] theoretically investigate the existence of Barren Plateaus for alternated disentangled UCC (dUCC) ansatz. Besides, Liu et al. [25] provide critical insights into the trainability of tensor-network models and reveal that barren plateaus are absent near the local minimum. This discovery has led to many investigations in new VQC architectures. For example, Martin et al. [26] investigate the BP issues in several randomly chosen parameterized circuits and reveal that some model architectures like tree tensor network (qTTN) and multi-scale entanglement re-normalization ansatz (qMERA) could avoid BPs. Cybulski et al. [27] investigate four scenarios about barren plateaus by examining the model depth, layer-by-layer pre-training, circuit block structure, and model creation without any constraints. Beyond evaluating VQCs, several works validate that BPs don’t exhibit in classical-quantum hybrid models. Abbas et al. [28] first investigate the advantages of quantum neural networks (QNNs) over classical neural networks and prove that certain QNNs can avoid BPs because of the unique optimization landscapes. Similarly, Pesah et al. [29] analyze the cost function of QCNNs and show that BPs don’t exhibit in QCNNs, indicating the potential of QCNNs for advancing NISQ computing. Coelho et al. [30] explore the trainability and performance of VQCs on deep Q-learning models in classical environments and indicate that increasing the number of qubits does not lead to BPs for these models.
@@ III-C Investigations from Optimization Perspectives
Another portion of the research works aims to investigate BPs from the perspective of the optimization process. Zhao et al. [31] propose a scheme to analyze the barren plateau phenomenon in training VQCs with the ZX-calculus. Cerezo et al. [8] study the connections between locality and trainability of different cost functions in shallow VQCs and reveal that optimizing VQCs via cost functions with global observables exhibit BPs, whereas the gradient vanishes polynomially when using local observables. Concurrently, Cerezo et al. [32] verify that Hessian-based approaches do not circumvent the exponential scaling associated with BPs. Later, Cerezo et al. [33] also reveal the cost function can be classically simulated when the gradient variance vanishes. This result arises from a curse of dimensionality. Current methods essentially encode this problem into small and classically simulatable subspaces. Arrasmith et al. [34] empirically verify that the gradient-free optimizer does not solve the barren plateau problem. Furthermore, Arrasmith et al. [35] investigate the interdependent connections among three optimization landscapes, barren plateaus, exponential cost concentration around the gradient mean, and narrow gorges, during the training phase of VQCs. This work establishes a connection between BPs and narrow gorges, implying that narrow gorges may occur under the BP landscape. Similarly, Miao et al. [36] evaluate BP issues and show that BPs would be addressed in the Riemannian formulation of such optimization problems. Furthermore, Miao et al. [37, 38] provide an in-depth analysis of isometric tensor network optimization methods applied to extensive Hamiltonians, and demonstrate that these approaches are inherently free of barren plateaus. Also, these works evaluate the gradient variance of tensor-network-based approaches and their scaling property. Pérez-Salinas et al. [39] investigate landscapes that contain relevant information causing optimization hardness, contributing connections between information content and the scale of the gradient. Nemkov et al. [40] investigate BPs in the presence of local traps and analyze how these traps deteriorate the difficulty of navigating barren plateaus, further hindering the optimization process. From the theoretical perspectives in optimization, Uvarov et al. [9] provide a lower bound for training VQCs and argue that the magnitude of gradient variance can be heavily influenced by the Hamiltonian and circuit structure. Letcher et al. [41] study the bounds of gradient descent in VQCs, providing tight upper and lower bounds during VQCs’ training for different amounts of observables.
@@ IV-A Initialization-based Strategies
Figure 2: A general structure of a VQC where the initialization-based strategy is circled in a dotted box. Generally speaking, initialization-based strategies $V(\theta)$ are applied to initialize model parameters $\theta$ of the unitary quantum circuits $U(\theta)$ before training.
@@
Grant et al. [46] propose a strategy that first randomly initializes some gates and further sets the rest gates to produce the identity matrix, ensuring initial gradients non-vanishing. Sauvage et al. [47] introduce a flexible initialization strategy that can adapt suitable initial parameters to various sizes of quantum circuits. Sack et al. [48] design a method to circumvent weak barren plateaus by using classical shadows protocol at the initialization stage and during the training process. Rad et al. [49] introduce a fast-and-slow algorithm that employs the Bayesian Learning approach to initialize parameters for VQCs by initializing a fast local optimizer to find the global optimum point efficiently. In this algorithm, the Bayesian approach can help predict effective starting points, thereby reducing the risk of barren plateaus. Kulshrestha et al. [50] first initialize the model parameters using a beta distribution, which can better model the data distribution in binary classification, and further add noise into the parameters during training to avoid the optimization being trapped in saddle points. They empirically verify that utilizing beta distribution can effectively mitigate the rapid decrease of gradient variance in initialization as the model size increases. Similarly, Zhang et al. [51] employ a Gaussian distribution as initial model parameters and demonstrate that such a strategy can effectively help VQCs avoid BPs. Friedrich et al. [52] utilize a classical neural network, namely a convolutional neural network (CNN), to randomly generate the model parameters, indicating the effectiveness of this method for mitigating BPs. Mele et al. [53] offer an iterative search scheme to transfer smooth solutions from smaller systems as a warm start for larger systems. Such a scheme has been proven to help optimization avoid the barren plateau issue. Grimsley et al. [54] first examine how the Adaptive Problem-Tailored Variational Quantum Eiegensolver (ADAPT-VQE) ansätze is affected by local minima and further propose a strategy that can initialize parameters with over an order of magnitude smaller error than random initialization. Liu et al. [55] present a transfer-learning approach where model parameters learned from small-scale tasks are transferred to larger tasks. This method leverages the smooth solutions from simpler problems to provide a good starting point for more complex VQCs, therefore avoiding barren plateaus and enhancing training efficiency. Park et al. [56] first show that a time-evolution operator generated by a local Hamiltonian can help VQCs avoid BPs and then propose an initialization scheme using this operator. Overall, initialization-based strategies can be summarized in Table II.

### v2

@@ Abstract
In recent years, variational quantum circuits (VQCs) have been widely explored to advance quantum circuits against classic models on various domains, such as quantum chemistry and quantum machine learning. Similar to classic machine-learning models, VQCs can be trained through various optimization approaches, such as gradient-based or gradient-free methods. However, when employing gradient-based methods, the gradient variance of VQCs may dramatically vanish as the number of qubits or layers increases. This issue, a.k.a. Barren Plateaus (BPs), seriously hinders the scaling of VQCs on large datasets. To mitigate the barren plateaus, extensive efforts have been devoted to tackling this issue through diverse strategies. In this survey, we conduct a systematic literature review of recent works from both investigation and mitigation perspectives. Furthermore, we propose a new taxonomy to categorize most existing mitigation strategies into five groups and introduce them in detail. Also, we compare the concurrent survey papers about BPs. Finally, we provide insightful discussion on future directions for BPs.
@@
Quantum Machine Learning, Variational Quantum Circuits, Mitigation of Barren Plateau, Survey
@@ I Introduction
In the era of noisy intermediate-scale quantum (NISQ), quantum computing has achieved significant advancement in various domains [1], such as quantum chemistry [2], quantum machine learning [3, 4, 5, 6], and quantum architecture [7, 8]. Despite these advancements, the superiority of quantum computing over classical computing in the foreseeable future remains uncertain. One uncertainty arises during the optimization of Variational Quantum Circuits (VQCs), which have emerged as pivotal models for advancing quantum computing. Similar to classical models, VQCs can be optimized by gradient-based approaches [9, 10]. During optimization, however, the training of VQCs may be initially trapped in a flat landscape as the model size increases. McClean et al. [11] first categorize this issue as barren plateaus (BPs) in VQCs and ascertain that the variance of the gradient will exponentially decrease as the model size increases when the VQCs match the 2-design Haar distribution. Under this occurrence, most gradient-based techniques would fail. Thus, it is essential to mitigate the barren plateau issues in VQCs.
@@
  • We conduct comprehensive literature reviews on recent works about barren plateaus (BPs) [^1] Our repo: https://github.com/junzhuang-code/Barren_Plateaus_Survey/.
@@ II Preliminary of Barren Plateaus
$U(\theta)=U(\theta_{1},...,\theta_{L})=\prod_{l=1}^{L}U_{l}(\theta_{l}),$ (1)
where $U_{l}(\theta_{l})=e^{-i\theta_{l}V_{l}}$, $V_{l}$ is a Hermitian operator, and $L$ denotes the number of layers in VQCs.
Similar to classical models, VQCs’ parameters can be optimized by gradient-based methods. The cost function for optimization can be denoted as $C(\theta)$. Thus, $C(\theta)$ can be formulated as the expectation over Hermitian operator $H$:
@@
Based on the above loss function $C(\theta)$, we can further denote its gradient as $\frac{\partial C(\theta)}{\partial\theta_{l}}$, a.k.a. $\partial C$. Note that, in this work, we use the term “optimization” to describe a process that searches optimal parameters for VQCs via numerical methods.
In 2018, McClean et al. [11] first investigated barren plateaus and revealed that under the assumption of the two-design Haar distribution, the variance of the gradient $Var[\partial C]$ in VQCs would exponentially decrease to zero during optimization as the model size, such as the number of qubits or layers, increases. This phenomenon is rooted in the assumption of the Haar distribution, under which $U(\theta)$ is considered sufficiently random in Hilbert space, making the gradient landscape increasingly flat as the size of VQCs grows. This randomness, a.k.a. Haar randomness, is characterized by the left- and right invariance of the Haar measure $\mu(U)$ on the unitary group $U(N)$, which ensures that the distribution of any function $f(U)$ remains unchanged under transformations by $V$ as follows.
[…]
The invariance presented in Equation 3 plays a crucial role in ensuring the uniformity of random unitaries. To achieve such invariance, however, VQCs require large computational resources, which is impractical in both simulation and real-world scenarios. To mitigate this drawback, researchers often leverage unitary $t$-designs [13] for restricted classes of $f(U)$, which significantly reduce computational demands while approximating the properties of the Haar measure. Formally, suppose that we sample unitary $V_{i}$ with probability $p_{i}$, then a unitary ensemble $\{p_{i},V_{i}\}$ is a $t$-design if it satisfies the following condition:
[…]
where $\rho$ denotes the density matrix, and $\otimes$ denotes the tensor product.
Equation 4 ensures that a $t$-design can replicate the statistical properties of Haar randomness for functions $f(U)$ of degree $t$ or lower. In practice, this means that if $f(U)$ is a polynomial with a maximum degree of $t$ in unitary $U$, averaging $f(U)$ over the $t$-design $\{p_{i},V_{i}\}$ approximates the result of averaging it over the unitary group w.r.t. the Haar measure [11].
In brief, while the Haar measure represents the theoretical upper bound of VQCs’ expressivity, it is often unattainable in practice due to excessive computational resource demands. $t$-designs provide a practical alternative to enhance VQCs’ expressivity by approximating Haar randomness with significantly reduced costs.
While $t$-designs mitigate the computational challenges associated with Haar randomness, they can still result in highly random unitary operators, particularly as the degree increases. This introduces a critical issue in VQCs’ training known as barren plateaus (BPs). This issue poses significant challenges to the practical deployment of VQCs. To generally define the BPs, Qi et al. [12] provide a simple but generic definition as follows:
@@
where $Var[\partial C]$ denotes the variance of gradient, $F(N)\in o\left(\frac{1}{b^{N}}\right)$, for some $b>1$ and, $N$ denotes the number of qubits in VQCs.
According to Equation 5, $Var[\partial C]$ will exponentially decrease to zero when the number of qubits $N$ increases. In this case, most gradient-based approaches will fail to train VQCs. Therefore, it is critical to mitigate barren plateau issues to train VQCs robustly. Notably, in this survey, we aim to investigate the mitigation methods designed to avoid the occurrence of barren plateaus at the beginning of the training process. Most mitigation methods theoretically or empirically verify the effectiveness under certain conditions, such as specific ranges of qubits or layers. The results indicate that the proposed methods can avoid BPs under these circumstances but do not necessarily guarantee the elimination of BPs for all cases. Thus, we use the term “mitigate” to describe these efforts.
@@ III-A Understanding Barren Plateau Phenomena
To dive deeper into BPs, Wang et al. [14] further prove that the gradient will vanish exponentially under the consideration of local Pauli noise, which is quite different from the noise-free setting in [11]. Similarly, Singkanipa et al. [15] explore the effects of noise on VQCs, focusing on how non-unital noise contributes to the formation of barren plateaus and fixed points in the optimization landscape. Marrero et al. [16] show that excessive entanglement between the visible and hidden units in VQCs can hinder learning capacity. Holmes et al. [17] first investigate the scrambling processes and show that showing that any variational ansatz is highly probable to have a barren plateau landscape. Later, Holmes et al. [18] study the expressibility of VQCs and reveal that highly expressive VQCs may exhibit flatter optimization landscapes. Similarly, Ragone et al. [19] provide a unified framework to study the expressibility of VQCs. Liu et al. [20] theoretically provide an understanding of the distinction between BPs and “laziness”. In this study, laziness refers to the exponential suppression of variational angle updates during gradient descent. By studying laziness, researchers have a deeper understanding of the ansatz structure as a trade-off between flat landscapes and trainability, i.e., trainability in the presence of BPs via laziness). Friedrich et al. [21] examine how the dimensionality of qudits influences the occurrence of BPs in VQCs. Fontana et al. [22] introduce the concept of adjoint differentiation to analyze BPs and provide theoretical insights for designing quantum ansätze to avoid barren plateaus. Comparably, Diaz et al. [23] extend the theoretical understanding of BPs beyond the conventional dynamical Lie algebra (DLA) framework and propose new criteria to identify BPs, while Goh et al. [24] leverage Lie-algebraic tools [25] to develop a classical simulation method, g-sim, that efficiently simulates BP-free unitary circuits. Last, Anschuetz et al. [26] investigate other untrainable scenarios beyond barren plateaus.
@@ III-B Investigations from Model Perspectives
A portion of the research works aims to investigate BPs from the perspective of existing quantum circuits. Wiersema et al. [27] first examine a family of VQCs, named the Hamiltonian Variational Ansatz (HVA), and reveal that HVA is absent from BPs due to its favorable structural properties. Mao et al. [28] theoretically investigate the existence of Barren Plateaus for alternated disentangled UCC (dUCC) ansatz. Besides, Liu et al. [29] provide critical insights into the trainability of tensor-network models and reveal that barren plateaus are absent near the local minimum. This discovery has led to many investigations in new VQC architectures. For example, Martin et al. [30] investigate the BP issues in several randomly chosen parameterized circuits and reveal that some model architectures like tree tensor network (qTTN) and multi-scale entanglement re-normalization ansatz (qMERA) could avoid BPs. Cybulski et al. [31] investigate four scenarios about barren plateaus by examining the model depth, layer-by-layer pre-training, circuit block structure, and model creation without any constraints. Beyond evaluating VQCs, several works validate that BPs don’t exhibit in classical-quantum hybrid models. Abbas et al. [32] empirically demonstrate that a class of quantum neural networks (QNNs) can outperform comparable feedforward networks and prove that certain QNNs can avoid BPs because of the unique optimization landscapes. Similarly, Pesah et al. [33] analyze the cost function of QCNNs and show that BPs don’t exhibit in QCNNs, indicating the potential of QCNNs for advancing NISQ computing. Coelho et al. [34] explore the trainability and performance of VQCs on deep Q-learning models in classical environments and indicate that increasing the number of qubits does not lead to BPs for these models.
@@ III-C Investigations from Optimization Perspectives
Another portion of the research works aims to investigate BPs from the perspective of the optimization process. Zhao et al. [35] propose a scheme to analyze the barren plateau phenomenon in training VQCs with the ZX-calculus. Cerezo et al. [9] study the connections between locality and trainability of different cost functions in shallow VQCs and reveal that optimizing VQCs via cost functions with global observables exhibit BPs, whereas the gradient vanishes polynomially when using local observables. Concurrently, Cerezo et al. [36] verify that Hessian-based approaches do not circumvent the exponential scaling associated with BPs. Later, Cerezo et al. [37] also reveal the cost function can be classically simulated when the gradient variance vanishes. This result arises from a curse of dimensionality. Current methods essentially encode this problem into small and classically simulatable subspaces. Arrasmith et al. [38] empirically verify that the gradient-free optimizer does not solve the barren plateau problem. Furthermore, Arrasmith et al. [39] investigate the interdependent connections among three optimization landscapes, barren plateaus, exponential cost concentration around the gradient mean, and narrow gorges, during the training phase of VQCs. This work indicates that in high-dimensional space, BPs will cause cost concentration and thus reduce the training efficiency. Similarly, Miao et al. [40] evaluate BP issues and show that BPs would be addressed in the Riemannian formulation of such optimization problems. Furthermore, Miao et al. [41, 42] provide an in-depth analysis of isometric tensor network optimization methods applied to extensive Hamiltonians, i.e., VQCs with extensive size, and demonstrate that these approaches are inherently free of barren plateaus. Also, these works evaluate the gradient variance of tensor-network-based approaches and their scaling property. Cao et al. [43] explore how many-body localization (MBL)-thermalization phase transitions can be utilized to avoid BPs, indicating that MBL states can be leveraged to optimize the design of quantum circuits. Pérez-Salinas et al. [44] investigate landscapes that contain relevant information causing optimization hardness, contributing connections between information content and the scale of the gradient. Nemkov et al. [45] investigate BPs in the presence of local minima and reveal that enlarging gradients cannot sufficiently address BPs under the above circumstance. From the theoretical perspectives in optimization, Uvarov et al. [10] provide a lower bound for training VQCs and argue that the magnitude of gradient variance can be heavily influenced by the Hamiltonian and circuit structure. Letcher et al. [46] study the bounds of gradient descent in VQCs, providing tight upper and lower bounds during VQCs’ training for different amounts of observables.
@@ IV-A Initialization-based Strategies
Figure 2: A general structure of a VQC where the initialization-based strategy is circled in a dotted box. Generally speaking, a VQC will take the encoded quantum-state data as inputs and learn the data representation through a sequence of unitary operations. For each unitary gate, VQC has a corresponding parameter. In the beginning, initialization-based strategies $V(\theta)$ are applied to initialize parameters $\theta$ of the unitary quantum circuits $U(\theta)$ before training. In the end, a measurement layer is applied to measure the output of the VQC.
@@
Grant et al. [50] propose a strategy that first randomly initializes some gates and further sets the rest gates to produce the identity matrix, ensuring initial gradients non-vanishing. Sauvage et al. [51] introduce a flexible initialization strategy that can adapt suitable initial parameters to various sizes of quantum circuits. Sack et al. [52] design a method to circumvent weak barren plateaus by using classical shadows protocol at the initialization stage and during the training process. Rad et al. [53] introduce a fast-and-slow algorithm that employs the Bayesian Learning approach to initialize parameters for VQCs by initializing a fast local optimizer to find the global optimum point efficiently. In this algorithm, the Bayesian approach can help predict effective starting points, thereby reducing the risk of barren plateaus. Kulshrestha et al. [54] first initialize the model parameters using a beta distribution, which can better model the data distribution in binary classification, and further add noise into the parameters during training to avoid the optimization being trapped in saddle points. They empirically verify that utilizing beta distribution can effectively mitigate the rapid decrease of gradient variance in initialization as the model size increases. Similarly, Zhang et al. [55] employ a Gaussian distribution as initial model parameters and demonstrate that such a strategy can effectively help VQCs avoid BPs. Friedrich et al. [56] utilize a classical neural network, namely a convolutional neural network (CNN), to randomly generate the model parameters, indicating the effectiveness of this method for mitigating BPs. Mele et al. [57] offer an iterative search scheme to transfer smooth solutions from smaller systems as a warm start for larger systems. Such a scheme has been proven to help optimization avoid the barren plateau issue. Grimsley et al. [58] first examine how the Adaptive Problem-Tailored Variational Quantum Eiegensolver (ADAPT-VQE) ansätze is affected by local minima and further propose a strategy that can initialize parameters with over an order of magnitude smaller error than random initialization. Liu et al. [59] present a transfer-learning approach where model parameters learned from small-scale tasks are transferred to larger tasks. This method leverages the smooth solutions from simpler problems to provide a good starting point for more complex VQCs, therefore avoiding barren plateaus and enhancing training efficiency. Park et al. [60] first show that a time-evolution operator generated by a $k$-local Hamiltonian can help VQCs avoid BPs and then propose an initialization scheme using this operator. The $k$-local Hamiltonian can perform efficiently as it acts non-trivially on at most $k$ qubits [61, 62], such as addressing the contains with finite correlation length in quantum spin lattice systems [63]. Some studies discussed in this survey also utilize the concept of $k$-local Hamiltonian [52, 64]. Overall, initialization-based strategies can be summarized in Table II.